39. A hockey player is standing on his skates on a frozen pond when an opposing player, moving with a uniform speed of 18 m/s, skates by with the puck. After 2.0 s, the first player makes up his mind to chase his opponent. If he accelerates uniformly at 4.3 m/s2, (a) how long does it take him to catch his opponent, and (b) how far has he traveled in this time? (Assume the player with the puck remains in motion at constant speed.) ef 2.1

please help. it's my physics class.
Answer #1

here,

the uniform speed of player 1 , u1 = 18 m/s

t1 = 2 s

accelration of seccond player , a2 = 4.3 m/s^2

let the time taken to catch be t and distance travelled by seccond player be s

for first player

s = (u1 * t1 + u1 * t)

s = ( 18 * 2 + 18 * t) ....(1)

and

for seccond player

s = 0 + 0.5 * a2 * t^2

s = 0 + 0.5 * 4.3 * t^2 ....(2)

from (1) and (2)

s = 216.7 m

t = 10.04 m

a)

the time taken is 10.04 m

b)

the distance travelled is 216.7 m